OpenStudy (anonymous):

Find the inverse function for f(x)=cos(7x), with 0≤x≤π/14. Find the inverse function for f(x)=cos(7x), with 0≤x≤π/14. @Mathematics

OpenStudy (anonymous):

but there is a domain for x

OpenStudy (anonymous):

x = cos(7y) arccosx=7y y=(1/7)arccosx but 0<=x<=pi/14 so ..... lets see